# Moving Aggregate Modified Autoregressive Copula-Based Time Series Models (MAGMAR-Copulas)

### v2

@@ Abstract
Copula-based time series models implicitly assume a finite Markov order. In reality a time series may not follow the Markov property. We modify the copula-based time series models by introducing a moving aggregate (MAG) part into the model updating equation. The functional form of the MAG-part is given as the inverse of a conditional copula. The resulting MAG-modified Autoregressive Copula-Based Time Series model (MAGMAR-Copula) is discussed in detail and distributional properties are derived in a D-vine framework. The model nests the classical ARMA model and can be interpreted as a non-linear generalization of the ARMA-model. The modeling performance is evaluated by modeling US inflation. Our model is competitive with benchmark models in terms of information criteria.
@@ 1 Introduction
Copula-based time series models are a class of probabilistic autoregressive time series models which allow for flexible temporal dependence modeling as well as flexible marginal modeling. The idea amounts to modeling (possibly non-linear) autoregressive temporal dependencies by a copula while modeling marginal properties independently. Given a univariate, strictly stationary time series $\{X_{t}\}_{t=1}^{T}$, the model decomposes the joint distribution, $F_{X_{t},\ldots,X_{t-p}}$ of $(p+1)$ consecutive random variables of the time series into their stationary marginal distribution, $F_{X}:=F_{X_{t}}=F_{X_{t-1}}=\ldots=F_{X_{t-p}}$, and their copula $C=F_{F_{X}(X_{t}),\ldots,F_{X}(X_{t-p})}$. The marginal distribution, $F_{X}$, accounts for marginal properties such as mean, variance, skewness etc., while the copula accounts for the temporal dependence structure. The general idea was formulated by Darsow et al. (1992) for Markov chains and further investigated in (amongst others) Ibragimov (2005). A short review is given in Patton (2009). Chen and Fan (2006) and Beare (2010) developed the idea for time series. Simard and Rémillard (2015) explore multivariate copula-based time series forecasting with standard copulas while Beare and Seo (2015), Brechmann and Czado (2015) and Nagler et al. (2022) explore multivariate time series modeling using vine-copulas.
@@
These models can be used to model non-linearities in temporal dependencies such as asymmetry; see Müller and Reuber (2022) for irradiation data and McNeil and Bladt (2022) for US inflation. For modeling of heavy-tailedness in the temporal dependence, see Loaiza-Maya et al. (2018) and Pappert and Arsova (2023). Linear temporal dependencies may also be modeled simply, by employing the normal copula. When choosing the copula to be the normal copula and the marginal distribution to be the normal distribution, the linear AR model is recovered (Chen and Fan, 2006, Smith et al., 2010). Thus the model can be understood as a generalization to the AR model which permits broader temporal dependencies and marginal distributions.
One caveat of the copula-based time series models is that, implicitly, the process is ascribed to be a Markov process of order $p$, i.e. the true data-generating process is of finite autoregressive order $p$. This means that the variable $X_{t}$ may depend explicitly on a finite number of past values $X_{t-1},\ldots,X_{t-p}$ only. However, in reality the Markov assumption may fail depending on the time series. The time series that is to be modeled may display persistent behavior i.e. non cut-off autodependence. In principle it is possible to extend the copula-based time series model to arbitrary Markov order, however it becomes unfeasible as the number of parameters increases indefinitely. Hence, we think that there is a need for a copula-based time series model with infinite order partial dependence. To this end, McNeil and Bladt (2022) propose a possible solution: They consider the copula-based time series model with an infinite autoregressive order. Hence, for each autoregressive order, there is a partial copula describing the corresponding temporal dependence. Handling such a model is, without further manipulation, impossible because it has infinitely many parameters. However, when reparametrizing it in a clever way, the model becomes manageable. The reparametrization proposed by the authors is based on linear ARMA models. The caveat of their model is that it is cumbersome to incorporate copulas with more than one parameter, e.g. the $t$-copula. In contrast, our model can accommodate general copulas by following a conceptually different approach. The approach by McNeil and Bladt (2022) is reviewed in Section 2.4.
Another approach to generalize copula-based time series models to allow for infinite autodependence is given in Chapter 3.14 of Joe (2014). Therein, a model, which is quite similar to the MAGMAR-copula model, is introduced. In particular, what we call ’moving aggregate’, he calls a ’$q$-dependent’ extension, and a copula extension to ARMA processes is named ’Mixed Markov and $q$-dependent’. Though similar, the models are not the same; while our model may be interpreted as feeding a non-linear AR model with noise filtered by a moving aggregate, the model by Joe feeds a non-linear AR-process to a non-linear $q$-dependent filter. We note that the MAGMAR-copula model has been developed independently. We discuss the similarities and differences between the two models in Sect. 2.5.
@@
Now we turn our attention to our proposed model. The problem of limited feasible autoregressive orders in time series models also occures in the AR($p$) and ARCH($p$) models. The solution in the AR($p$) and ARCH($p$) models is to introduce moving average terms, resulting the the ARMA and GARCH models, respectively. (See Engle (1982), Bollerslev (1986), Bollerslev et al. (1994), Francq and Zakoian (2019) and Hamilton (2020)). By inverting the corresponding linear filter, AR($\infty$), respectively ARCH($\infty$) representations can be derived. Hence the time series can be understood as having no Markov restriction. This results in the time series models being more persistent in the mean in the case of ARMA and more persistent in the variance in the case of GARCH. The latter can be interpreted as the model being able to capture volatility clusters (Bollerslev et al., 1994).
@@
Inspired by this approach, we modify the copula-based time series model by incorporating a moving aggregate (MAG) part in the updating equation. The MAG part is a non-linear function taking the current and the past $q$ innovations as arguments. Formally it is a mapping from $\mathbb{R}^{q+1}$ to $\mathbb{R}$. The functional form of the MAG-function is given by an inverse conditional copula. We call it ’moving aggregate’, because, similar to the moving average, it aggregates the current and past $q$ innovations but it is no linear aggregation, hence ’aggregate’ instead of ’average’. Analogous to the classical ARMA and GARCH models, the MAG-part permits infinite partial dependence. With this, we mean that the partial copula (see Spanhel and Kurz, 2016) of $X_{t}$ and $X_{t-j}$ is different from the independence copula for all $j,t\in\mathbb{Z}$. We call the resulting model MAG-Modified Autoregressive Copula-Based Time Series Model, or in short: MAGMAR-Copula. In this paper we present the model and show that the classical copula-based time series model and the classical linear ARMA model are nested within the proposed model. We proceed to derive its distributional properties in a D-vine framework (see Aas et al. (2009), Czado (2010), Czado and Nagler (2022) and Sect. 2.1 for D-vine copulas). AR($\infty$) and MAG($\infty$) representations are derived. Building upon the MAG($\infty$) representation, it is proven that the MAGMAR-Copula model, where the AR-copula is the normal copula, is stationary when excluding perfect comonotonicity and perfect countermonotonicity. Furthermore, we derive the conditional distributions of $X_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{t-1}$ denotes the $\sigma$-field of all information up to and including $(t-1)$. Based on conditional independence, the likelihood function is derived. Similar to ARMA and GARCH models, the likelihood has to be calculated in an iterative scheme.
Last, we apply the MAGMAR-Copula model to quarterly US inflation. The parameters are estimated by maximum likelihood estimation. We compare the modeling performance of the MAGMAR-Copula model to that of the copula-ARMA model of McNeil and Bladt (2022) and to linear modeling (where normal copulas as temporal dependence structure). We find that our model performs well in this setting and is competitive with the model from McNeil and Bladt (2022). In line with their results, we find that modeling asymmetric temporal dependence structures (by incorporating the gumbel copula) as well as infinite partial dependence is beneficial. Furthermore, we find that incorporating heavy-tailed temporal dependence structures (i.e. by incorporating the $t$-copula) enhances the performance, indicating not only asymmetric but also heavy-tailed temporal dependencies in US inflation.
@@
In the next section, Sect. 2, the copula-based time series model is introduced and reviewed. We focus on the model class where the copula is modeled as a D-Vine copula. Sect. 2.4 reviews the model for time series with infinite partial copula dependence by McNeil and Bladt (2022). In Sect. 3 the Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula) is introduced and explored. The MAGMAR-Copula model is applied to US Inflation in Sect. 4. Sect. 5 concludes the work and gives an outlook to future research.
@@ 2 Copula-based Time Series Models
In this section the notion of copula-based time series models is introduced. First, we give a short introduction to copulas. Then general ideas of copula-based time series is developed for the Markov(1) case and then extended to arbitrary order. Last, we discuss the copula-ARMA model, which permits infinite partial dependence, proposed by McNeil and Bladt (2022).
@@ 2.1 Copulas and D-vine Copulas
Since the copula density is given by the partial derivative of the respective copula, Eq. 2 is only valid if the copula is differentiable. In applications, the copula and the marginal distributions have to chosen appropriately. The normal, t, gumbel and independence copula are described in G. There are other important copulas that deserve being mentioned here, however, we restrict the exposition to the copulas that are used in this work for brevity. Again we refer to Joe (2014) and Nelsen (2006) for more examples of copulas.
@@
where $c_{i,i+j|i+1,\ldots,i+j-1}$ is the conditional copula density of $U_{i}$ and $U_{i+j}$ given the intermediate variables $U_{i+1},\ldots,U_{i+j-1}$. $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ is the conditional distribution function of $U_{i}$ given $U_{i+1},\ldots,U_{i+j-1}$ evaluated at $u_{i}$ and conditioning variables $u_{i+1},\ldots,u_{i+j-1}$. Since the joint distribution of $U_{i},U_{i+1},\ldots,U_{i+j-1}$ is a copula, $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ may also be understood as a conditional copula. The decomposition in Eq. 3 is called ’D-vine’ for ’drawable vine’ because the variables are ordered in a path like manner (Czado and Nagler, 2022). The vine copula approach allows for more flexibility when modeling multivariate dependencies because each bivariate dependence can be specified separately.
@@ 2.2 Copula-based Time Series with Autoregressive Order of One
Consider a strictly stationary time series $\{X_{t}\}_{t=1}^{T}$. A time series being Markov(1) means that $\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{1}=x_{1})=\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1}),\ t>1$. Accordingly a time series of Markov order $p$ fulfills $\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{1}=x_{1})=\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{t-p}=x_{t-p}),\ t>p$ (Ibe, 2013). First, we consider the case where $p=1$. Afterwards, we consider the case for arbitrary $p$. For $p=1$, the time series is Markov($1$) and characterized by the bivariate joint distribution $F_{X_{t},X_{t-1}}$ of consecutive random variables, $X_{t}$ and $X_{t-1}$. The idea is to decompose the joint distribution into a copula and marginal distributions. By Sklar’s Theorem there always exists a copula $C:=F_{F_{X}(X_{t}),F_{X}(X_{t-1})}$ such that the joint distribution can be decomposed into marginal distributions $F_{X_{t}}$ and $F_{X_{t-1}}$ and the copula (Sklar, 1959, Joe, 2014). If the random variables $X_{t}$ and $X_{t-1}$ are continuous, the decomposition is unique. Due to strict stationarity, $F_{X_{t}}=F_{X_{t-1}}=:F_{X}$. The decomposition may be written as
$F_{X_{t},X_{t-1}}(a,b)=C[F_{X}(a),F_{X}(b)].$
@@
Hence, the joint distribution is specified by a marginal distribution $F_{X}$ and the copula $C$. Note that the copula is the joint distribution of the probability-integral transformed random variables $U_{t}:=F_{X}(X_{t})$ and $U_{t-1}:=F_{X}(X_{t-1})$. The copula may be parametrized by dependence parameter(s) $\phi$, so $C=C_{\phi}$. We introduce this notation to emphasize the analogy to the linear AR model and later on, in Sect. 3 to the linear ARMA model. Then, analogous to the linear AR model, the Markov($1$) model can be formulated in a time series model equation as
$X_{t}=F^{-1}_{X}\left(C_{\phi}^{-1}[w_{t}|F_{X}(X_{t-1})]\right),\ \ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$
@@
We note that this or similar formulations are also used in other works under different names such as ’updating equation’ in McNeil and Bladt (2022) or in descriptions on how to sample from copulas in Aas et al. (2009). The calculation of the new realization, $x_{t}$, amounts to first sampling a realization from a uniform distribution, $w_{t}$ and then quantile transforming $w_{t}$ with the conditional copula obtaining $U_{t}:=C_{\phi}^{-1}[w_{t}|F_{X}(X_{t-1})]$. Last, $U_{t}$ is again quantile transformed with the marginal distribution quantile function $F_{X}^{-1}$. Using this procedure, it is guaranteed that the dependence structure of $X_{t}$ and $X_{t-1}$ is given by $C_{\phi}$ and that the marginal distribution is given by $F_{X}$. This model can be regarded as a generalization of the gaussian AR(1)-model. This can be seen by choosing $C_{\phi}$ as the bivariate normal copula $C[u_{1},u_{2}]=\Phi_{\phi}(\varphi^{-1}(u_{1}),\varphi^{-1}(u_{2}))$ (see e.g. Joe (2014) for details about the normal copula or Sect. 2.1) and the marginal distribution $F_{X}$ as the standard normal distribution. Although this is a well known fact in the literature, see e.g. Chen and Fan (2006) or Smith et al. (2010), we give the derivation in Appendix A for completeness. In the following only the pseudo-observation time series, $\{U_{t}\}_{t=1}^{T},\ U_{t}:=F_{X}(X_{t})$, is considered. The actual time series can be recovered by quantile transforming every pseudo-obersvation by $F_{X}^{-1}$. The dynamics of the time series are solely determined by the relation between the pseudo-observations. The model equation of the pseudo-observations is given as,
$U_{t}=C^{-1}_{\phi}[w_{t}|U_{t-1}],\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (4)
The non-linearity of the model is two-fold. First, the relation between random variables $U_{t}$ and $U_{t-1}$ is a non-linear function. Second, the innovations $w_{t}$ enter in a non-additive way. What at first sight may seem strange is actually necessary. Choosing the model equation this way is the only possibility to ensure that $U_{t},U_{t-1},\ldots$ are uniformly distributed and that $C_{\phi}$ is the copula of consecutive variables. The uniformity of the unconditional distribution of all $U_{t},\ t\in\mathbb{Z}$ is due the properties of inverse conditional copulas: Given $U_{1},U_{2}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ it follows that $C^{-1}_{\phi}[U_{1}|U_{2}]\sim U(0,1)$. This implies that all $U_{t}$ are distributed identically. From this it follows that the time series is stationary, because the transitions in the updating equation, Eq. 4, are homogeneous.
2.3 Copula-based Time Series models with an arbitrary autoregressive order
The model can be generalized in the case where $\{X_{t}\}_{t\in\mathbb{Z}}$ is a strictly stationary Markov($p$) time series. The behavior of the time series is determined by the joint distribution of $p+1$ consecutive random variables $X_{t},X_{t-1},\ldots,X_{t-p}$. As before, decomposing the joint distribution using Sklar’s Theorem yields
$F_{X_{t},\ldots,X_{t-p}}(x_{t},\ldots,x_{t-p})=C[F_{X}(x_{t}),\ldots,F_{X}(x_{t-p})],$
where again $F_{X}:=F_{X_{t}}=\ldots=F_{X_{t-p}}$ due to strong stationarity and $C$ is the joint distribution of the individually uniformly distributed random variables $(U_{t},\ldots,U_{t-p})$ where $U_{t}:=F_{X}(X_{t})\ \forall t\in\mathbb{Z}$. Considering the time series of pseudo-observations $\{U_{t}\}_{t=1}^{T}$ and parametrizing the copula with dependence parameter(s) $\phi$, the time series model equation can be formulated as
$U_{t}=C_{\phi}^{-1}[w_{t}|U_{t-1},\ldots,U_{t-p}],\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (5)
@@ 2.3 Copula-based Time Series models with an arbitrary autoregressive order
Note that the copula $C_{\phi}$ is now a $(p+1)$-dimensional copula. The parameter $\phi$ may be a scalar or a vector, depending on the chosen copula. Choosing a specific copula as $C_{\phi}$ may restrict the model to one dependence structure. In order to account for possibly different forms of dependency and to be flexible in modeling, D-vine copulas (Bedford and Cooke, 2001, 2002, Aas et al., 2009, Czado, 2010, Czado and Nagler, 2022) can be employed for modeling the temporal dependence as is also done in Nagler et al. (2022) and Smith et al. (2010). In this approach the joint density of $(p+1)$ consecutive pseudo-observations is decomposed into a product of bivariate copulas. The decomposition is given by (Smith et al., 2010, McNeil and Bladt, 2022, Nagler et al., 2022)
@@
where the shorthand notation $c_{t,\ldots,t-p}=c_{U_{t},\ldots,U_{t-p}}$ is used. Furthermore, $S_{j-k,j}=\{j-k+1,\ldots,j-1\}$. As in the cited literature, only the simplified D-vine is considered, dropping the conditioning in the bivariate copula densities, $c_{j-k,j|S_{j-k,j}}=c_{j-k,j}$. In the case where $c_{j-k,j}=c_{i-k,i}:=c_{k}\ \forall j,i\in\{k+1,\ldots,T\}$, the joint density of all random variables $\{U_{t}\}_{t=1}^{T}$ admits translational invariance. Nagler et al. (2022) show that such processes are stationary. They also may be called s-vine (’s’ for stationary) (McNeil and Bladt, 2022). Furthermore, the bivariate copulas, $C_{j-k,j}$ may be parametrized by parameters $\phi_{1},\ldots,\phi_{p}$, where $\phi_{1}$ governs the dependence of $U_{t-1}$ and $U_{t}$ $\forall t\in\mathbb{Z}$, $\phi_{2}$ the dependence of $U_{t}|U_{t-1}$ and $U_{t-2}|U_{t-1}$ and so on. Note that depending on the chosen copula, each $\phi_{i},i\in\{1,\ldots,p\}$ may also be a vector containing more than one parameter.

### v3

@@ Abstract
Copula-based time series models can model univariate and stationary time series in a flexible way by decomposing the joint distribution of consecutive observations into a copula and the stationary distribution. Implicitly this approach assumes a finite Markov order. In reality a time series may not follow the Markov property. We modify the copula-based time series models by introducing a moving aggregate (MAG) part into the model updating equation. The functional form of the MAG-part is given as the conditional quantile function corresponding to a copula. The resulting MAG-modified Autoregressive Copula-Based Time Series model (MAGMAR-Copula) is discussed in detail and distributional properties are derived in a D-vine framework. We show that the stationary distribution implied by the model is not standard-uniform. Hence we propose an adjustment transformation that recovers the desired standard-uniformity. The model nests the classical ARMA model and can be interpreted as a non-linear generalization of the ARMA-model. The modeling performance is evaluated by modeling US inflation. Our model is competitive with benchmark models in terms of information criteria.
@@ 1 Introduction
Copula-based time series models are a class of univariate probabilistic autoregressive time series models which allow for flexible temporal dependence modeling as well as flexible marginal modeling. The idea amounts to modeling (possibly non-linear) autoregressive temporal dependencies by a copula while modeling marginal properties independently. Given a univariate, strictly stationary time series $\{X_{t}\}_{t\in\mathbb{Z}}$, the model decomposes the joint distribution, $F_{X_{t},\ldots,X_{t-p}}$ of $(p+1)$ consecutive random variables of the time series into their stationary marginal distribution, $F_{X}:=F_{X_{t}}=F_{X_{t-1}}=\ldots=F_{X_{t-p}}$, and their copula $C=F_{F_{X}(X_{t}),\ldots,F_{X}(X_{t-p})}$. The marginal distribution, $F_{X}$, accounts for marginal properties such as mean, variance, skewness etc., while the copula accounts for the temporal dependence structure. The general idea was formulated by Darsow et al. (1992) for Markov chains and further investigated in (amongst others) Ibragimov (2005). A short review is given in Patton (2009). Chen and Fan (2006) and Beare (2010) developed the idea for time series. Simard and Rémillard (2015) explore multivariate copula-based time series forecasting with standard copulas while Beare and Seo (2015), Brechmann and Czado (2015) and Nagler et al. (2022) explore multivariate time series modeling using vine-copulas.
@@
These models can be used to model non-linearities in temporal dependencies such as asymmetry; see Müller and Reuber (2022) for irradiation data and McNeil and Bladt (2022) for US inflation. For modeling of heavy-tailedness in the temporal dependence, see Loaiza-Maya et al. (2018) and Pappert and Arsova (2023). Linear temporal dependencies may also be modeled simply by employing the normal copula. Thus the model can be understood as a generalization to the AR model, which permits broader temporal dependencies and marginal distributions.
@@
One caveat of the copula-based time series models is that, implicitly, the process is ascribed to be a Markov process of order $p$, i.e. the true data-generating process is of finite autoregressive order $p$. This means that the variable $X_{t}$ may depend on a finite number of past values $X_{t-1},\ldots,X_{t-p}$ only. However, in reality the Markov assumption may fail depending on the time series. The time series that is to be modeled may display persistent behavior i.e. non cut-off autodependence. A popular example of this behavior is volatility clustering. There, persistency is apparent in the second moment of the time series. In principle it is possible to extend the copula-based time series model to arbitrary Markov order, however it becomes unfeasible as the number of parameters increases indefinitely. Hence, we think that there is a need for a copula-based time series model with infinite order partial dependence. To this end, we take inspiration from classic time series models. The problem of limited feasible autoregressive orders in time series models also occurs in the AR($p$) and ARCH($p$) models. The solution for these models is to introduce moving average terms, resulting in the ARMA(p,q)(p,q) and GARCH(p,q)(p,q) models, respectively (see Engle (1982), Bollerslev (1986), Bollerslev et al. (1994), Francq and Zakoian (2019) and Hamilton (2020)). By inverting the corresponding linear filter, AR($\infty$) and ARCH($\infty$) representations can be derived. Hence the time series can be understood as having no Markov restriction. This results in the time series models being more persistent in the mean in the case of ARMA and more persistent in the variance in the case of GARCH. The latter can be interpreted as the model being able to capture volatility clusters (Bollerslev et al., 1994). See also Fig. 1 and Fig. 2.
Figure 1: Simulations of an AR(3)(3) and a ARMA(3,1)(3,1) process using the same random number generation. The ARMA process exhibits long-term autoregressive effects in the mean, i.e. smoothness.
Figure 2: Simulations of an ARCH(1)(1) and a GARCH(1,1)(1,1) process using the same random number generation. The GARCH process exhibits long-term autoregressive effects in the variance, i.e. volatility clustering.
@@
Inspired by these approaches, we modify the copula-based time series model by incorporating a moving aggregate (MAG) part in the updating equation. This allows for long-term autoregressive effects in the whole conditional distribution. The MAG part is a non-linear function taking the current and the past $q$ innovations as arguments. Formally it is a mapping from $(0,1)^{q+1}$ to $(0,1)$. The functional form of the MAG-function is given by a conditional quantile function corresponding to a copula. We call it ’moving aggregate’, because, similar to the moving average, it aggregates the current and past $q$ innovations but it is not a linear aggregation, hence ’aggregate’ instead of ’average’. Analogous to the classical ARMA and GARCH models, the MAG-part permits infinite partial dependence. This means that the partial copula (see Spanhel and Kurz, 2016) of $X_{t}$ and $X_{t-j}$ is in general different from the independence copula for all $j,t\in\mathbb{Z}$. We call the resulting model MAG-Modified Autoregressive Copula-Based Time Series Model, or in short: MAGMAR-Copula. In this paper we present the model and show that the classical copula-based time series model and the classical linear ARMA model are nested within the proposed model. We proceed to derive its distributional properties in a D-vine framework (see Aas et al. (2009), Czado (2010), Czado and Nagler (2022) and Sect. 2.1 for D-vine copulas). AR($\infty$) and MAG($\infty$) representations are derived. Building upon the MAG($\infty$) representation, it is first proven that the MAGMAR(1,1)(1,1)-Copula model, where the AR-copula is the normal copula, is stationary when excluding perfect comonotonicity and perfect countermonotonicity. Second, we establish stationarity and ergodicity under a contraction-on-average property of the time series. In general the stationary distribution of the time series is not uniform (as is desired). We introduce an auxiliary adjustment transformation to account for this problem. The transformation guarantees that the unconditional distribution of the time series can always be modeled properly. We derive the conditional distribution of $X_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{t-1}$ denotes the $\sigma$-field of all information up to and including (t1)(t-1). Based on conditional independence, the likelihood function is derived. Similar to ARMA and GARCH models, the likelihood has to be calculated in an iterative scheme. The algorithm is presented and asymptotic properties of the estimation are derived. Since asymptotic results for the estimation of Markov copula-based time series models are available in the literature (see Chen and Fan (2006) or Nagler et al. (2022)), we focus on the consistency of the maximum-likelihood estimation of the MAG(1)(1)-copula model. We proceed to systematically compare our proposed model with the generalizations of copula-based time series by McNeil and Bladt (2022) and Joe (2014) (chapter 3.14). These models also aim to overcome the Markov restriction/generalize the linear ARMA model. Last, we test the model in a simulation study and apply it to quarterly US inflation. We compare the modeling performance of the MAGMAR-Copula model to the model by McNeil and Bladt (2022) and to linear modeling (i.e. choosing normal copulas for the temporal dependence structure). We find that our model performs well in this setting and is competitive with the model from McNeil and Bladt (2022). In line with their results, we find that proper modeling of asymmetric temporal tail dependence as well as infinite partial dependence is beneficial. Furthermore, we find that incorporating heavy-tailed temporal dependence (i.e. incorporating the $t$-copula) enhances the performance, indicating not only asymmetric tail-dependence but also heavy-tailed dependencies in US inflation.
@@
In the next section, Sect. 2, the copula-based time series model is introduced and reviewed with a focus on D-vines. Furthermore related generalizations of copula-based time series models are reviewed. In Sect. 3 the Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula) is introduced and theoretical properties are derived. In Sect. 4, we explore the model in a simulation study, verifying theoretical findings, extending them and comparing the model with related ones. The MAGMAR-Copula model is applied to US Inflation in Sect. 5. Sect. 6 concludes the work and gives an outlook to future research.
@@ 2 Copula-based Time Series Models
In this section the notion of copula-based time series models is introduced. First, we give a short introduction to copulas. Then, general ideas of copula-based time series modeling are developed. Last, we discuss the approaches to generalize copula-based time series modeling by McNeil and Bladt (2022) and Joe (2014).
@@ 2.1 Copulas and D-vine Copulas
Since the copula density is given by the partial derivative of the respective copula, Eq. 2 is only valid if the copula is differentiable. In applications, the copula and the marginal distributions have to be chosen appropriately. The normal, $t$, Gumbel and independence copulas are described in C.1.
@@
where $c_{i,i+j|i+1,\ldots,i+j-1}$ is the conditional copula density of $U_{i}$ and $U_{i+j}$ given the intermediate variables $U_{i+1},\ldots,U_{i+j-1}$. $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ is the conditional distribution function of $U_{i}$ given $U_{i+1},\ldots,U_{i+j-1}$ evaluated at $u_{i}$ and conditioning variables $u_{i+1},\ldots,u_{i+j-1}$. Since the joint distribution of $U_{i},U_{i+1},\ldots,U_{i+j-1}$ is a copula, $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ may also be understood as a conditional distribution function, corresponding to a copula. The decomposition in Eq. 3 is called ’D-vine’ for ’drawable vine’ because the variables are ordered in a path like manner (Czado and Nagler, 2022). The vine copula approach allows for more flexibility when modeling multivariate dependencies because each bivariate dependence can be specified separately.
@@ 2.2 Copula-based Time Series with Autoregressive Order of One
Consider a strictly stationary time series $\{X_{t}\}_{t\in\mathbb{Z}}$. A time series being Markov(p)(p) means that $\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{1}=x_{1})=\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{t-p}=x_{t-p}),\ t>p$ (Ibe, 2013). A stationary and univariate Markov(p)(p) time series is characterized by the $(p+1)$-variate joint distribution $F_{X_{t},\ldots,X_{t-p}}$ of consecutive random variables $X_{t}$ to $X_{t-p}$. The idea is to decompose the joint distribution into a copula and marginal distributions. By Sklar’s Theorem there always exists a copula $C:=F_{F_{X}(X_{t}),\ldots,F_{X}(X_{t-p})}$ such that the joint distribution can be decomposed into marginal distributions $F_{X_{t}}$ to $F_{X_{t-p}}$ and the copula (Sklar, 1959, Joe, 2014). If the random variables $\{X_{t}\}$ are continuous, the decomposition is unique. Due to strict stationarity we can write $F_{X_{t}}=\ldots=F_{X_{t-p}}=:F_{X}$. Then the decomposition may be written as
$F_{X_{t},\ldots,X_{t-p}}(x_{0},\ldots,x_{p})=C[F_{X}(x_{0}),\ldots,F_{X}(x_{p})].$
@@
Hence, the joint distribution is specified by a marginal distribution $F_{X}$ and the copula $C$. Note that the copula is the joint distribution of the probability-integral transformed random variables $U_{t}:=F_{X}(X_{t})$ to $U_{t-p}:=F_{X}(X_{t-p})$. The copula may be parametrized by dependence parameter(s) $\phi$, so we write $C=C_{\phi}$ . We introduce this notation to emphasize the analogy to the linear AR model and later on, in Sect. 3 to the linear ARMA model. Then, analogous to the linear AR model, the Markov($p$) model can be formulated in terms of a updating equation as
$X_{t}=F^{-1}_{X}\left(C_{\phi}^{-1}[w_{t}|F_{X}(X_{t-1}),\ldots,F_{X}(X_{t-p})]\right),\ \ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$
@@
where $C_{\phi}^{-1}$ is the conditional quantile function corresponding to the copula $C_{\phi}$. Formally $C_{\phi}^{-1}[x|u_{1},\ldots,u_{p}]$ is the solution to $C_{\phi}[x|u_{1},\ldots,u_{p}]=x$. We note that this or similar formulations are also used in other works under different names such as ’updating equation’ in McNeil and Bladt (2022) or in descriptions on how to sample from copulas in Aas et al. (2009). This model can be regarded as a generalization of the Gaussian AR(p)(p)-model. This can be seen by choosing $C_{\phi}$ as the normal copula and the marginal distribution $F_{X}$ as the standard normal distribution. Although this is a well-known fact in the literature, see e.g. Chen and Fan (2006) or Smith et al. (2010), we give the derivation for $p=1$ in B.1 for completeness. In the following only the pseudo-observation time series, $\{U_{t}\}_{t=1}^{T},\ U_{t}:=F_{X}(X_{t})$, is considered. The actual time series can be recovered by quantile transforming every pseudo-observation by $F_{X}^{-1}$. The dynamics of the time series are solely determined by the relation between the pseudo-observations. The model equation of the pseudo-observations is given as,
$U_{t}=C^{-1}_{\phi}[w_{t}|U_{t-1},\ldots,U_{t-p}],\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (4)
@@
In order to account for possibly different forms of dependency and to be flexible in modeling, D-vine copulas (Bedford and Cooke, 2001, 2002, Aas et al., 2009, Czado, 2010, Czado and Nagler, 2022) can be employed for modeling the temporal dependence as is also done in Nagler et al. (2022) and Smith et al. (2010). In this approach the joint density of $(p+1)$ consecutive pseudo-observations is decomposed into a product of bivariate copulas. The decomposition is given by (Smith et al., 2010, McNeil and Bladt, 2022, Nagler et al., 2022)
@@
where the shorthand notation $c_{t,\ldots,t-p}=c_{U_{t},\ldots,U_{t-p}}$ is used. Furthermore, $S_{j-k,j}=\{j-k+1,\ldots,j-1\}$. As in the cited literature, only the simplified D-vine is considered, dropping the conditioning in the bivariate copula densities, $c_{j-k,j|S_{j-k,j}}=c_{j-k,j}$. In the case where $c_{j-k,j}=c_{i-k,i}:=c_{k}\ \forall j,i\in\{k+1,\ldots,T\}$, the joint density of all random variables $\{U_{t}\}_{t\in\mathbb{Z}}$ admits translational invariance. Nagler et al. (2022) show that such processes are stationary. Hence we refer to such D-vine copulas as stationary D-vine copulas. They also may be called s-vine (’s’ for stationary) (McNeil and Bladt, 2022). Furthermore, the bivariate copulas, $C_{j-k,j}$ may be parametrized by parameters $\phi_{1},\ldots,\phi_{p}$, where $\phi_{1}$ governs the dependence of $U_{t-1}$ and $U_{t}$ $\forall t\in\mathbb{Z}$, $\phi_{2}$ the dependence of $U_{t}|U_{t-1}$ and $U_{t-2}|U_{t-1}$ and so on. Note that depending on the chosen copula, each $\phi_{i},i\in\{1,\ldots,p\}$ may also be a vector containing more than one parameter.